# A phase transition in Barak-Erdős random graphs

### v1

@@ Abstract
We study monotone paths in Erdős-Rényi random graphs on numbered vertices. This model appeared in [Benjamini2022] where Benjamini & Tzalik established a phase transition at $p=\frac{\log n}{n}$. We refine the critical value to $p=\frac{\log n-\log\log n}{n}$ and identify the critical window of order $\Theta(1/n)$.
@@ 1 Introduction
Recall the definition of a Barak-Erdős random graph $G_{p}$ on the positive integers $\mathbb{N}$: A directed edge $(i,j)$ between any two vertices $i<j\in\mathbb{N}$ exists independently with probability $p\in(0,1)$. We denote this law by $\mathbb{P}_{p}$. We say that there is a monotone path from 11 to $n$, write $1\nearrow n$, if there exists a path $1=i_{1}\to i_{2}\to\cdots\to i_{k}=n$ in $G_{p}$ whith $i_{1}<i_{2}<\cdots<i_{k}$. We establish a phase transition for the increasing event
@@
Erdős-Rényi graphs on numbered vertices $\{1,\ldots,n\}$ appeared as an ingredient to recover point configurations in [Benjamini2022]. They prove a sharp threshold at $p=\frac{\log n}{n}$ for the event $1\nearrow n$. Directed acyclic Erdös-Rényi graphs were introduced in [barak.erdos1984] and studied as Barak-Erdős graphs since. The articles [foss2003extended], [mallein2021] and [FoKoMaRa2024] established maximal path lengths in Barak-Erdős graphs, both for fixed $p\in(0,1)$ and $p\to 0$.
Numbered vertices can be interpreted as time stamps: each vertex becomes available from its assigned time onward. Related models include random temporal graphs, where monotonicity of paths is considered along edges instead of vertices, see for instance [AngelFerber...2018], [CasteigtsRaskin2024] and [BroutinKamcevLugosi2024]. These works focus on the length of monotone paths and identify several transition phenomena around $p=c\frac{\log n}{n}$ for both longest and shortest such paths.
A vertex order appears naturally if the nodes correspond to events unfolding in time, as for example in graphical models of causality [pearl2009causality]. Erdős-Rényi graphs directed according to a random vertex order are of particular interest to research in causal discovery [spirtes2001causation], where they are widely used in the benchmarking of algorithms that learn causal graphs from data.
@@ 2 Proof
The central ingredient to prove Theorem 1.1 is the introduction of an exploration process. It indicates the position of having passed a certain number of vertices that are reachable from 11 by a monotone path. We show that, after suitable centering and scaling, this exploration process converges to a Gumbel variable with a deterministic drift in Section 2.1. This allows to compute explicitly the probability that a monotone path $1\nearrow n$ exists.
@@
the number of vertices up to $k$ that are reachable from 11 by a monotone path and set $\mathcal{F}_{k}=\sigma(N_{1},\ldots,N_{k})$. A new vertex $k$ is reachable from 11 precisely when there is an edge to at least one of the $N_{k-1}$ previously reachable vertices. Hence
@@
where each large bullet marks a vertex reachable from 11 and each group of small dots represents the vertices in the corresponding gap: The $i$-th gap is a geometric variable with parameter $1-(1-p)^{i}$, and the gaps are pairwise independent.
@@ 2.1 Exploration process
denotes the position of the $\lfloor a/p\rfloor$-th vertex reachable from 11 by a monotone path. Since $P_{p}(a)$ is obtained by summing the successive gap lengths, we have
@@
Define $q_{i}\coloneq 1-(1-p)^{i}$. Assume without loss of generality that $p<1/2$. For any $p<1/2$ and $i\leq\lfloor a/p\rfloor$ it holds $q_{i}\leq 1-(1-p)^{\lfloor a/p\rfloor}\leq B(a)<1$ for some constant $B(a)$ independent of pp. Each $P_{p}(a)$ is the sum of independent geometric variables with expectation $1/q_{i}$.
@@
so that $Y_{i}$ are exponential variables with intensity 11 and $X_{i}$ are geometric variables with mean $1/q_{i}$. Then $\log\left(\frac{1}{1-q_{i}}\right)(X_{i}-1)\leq Y_{i}$ which allows the bound
@@
We now conclude the convergence as a process in $a>0$ via a probabilistic version of Dini’s lemma, see [Curien2025, Theorem 11.6]: The process
@@
It remains to prove Lemma $2.$
Proof of Lemma $2.$
@@ 2.2 Critical window
Proof of Theorem $1.$
For each $n\in\mathbb{N}$ define the random variable $\mathfrak{X}_{n}$ as the number of vertices smaller than $n$ that are reachable from 11 by a monotone path, scaled by $p_{n}$, that is

### v2

@@ Abstract
We study monotone paths in Erdős-Rényi random graphs on numbered vertices. This model appeared in [3] where Benjamini & Tzalik established a phase transition at $p=\frac{\log n}{n}$. We refine the critical value to $p=\frac{\log n-\log\log n}{n}$ and identify the critical window of order $\Theta(1/n)$.
@@ 1 Introduction
Recall the definition of a Barak-Erdős random graph $G_{p}$ on the positive integers $\mathbb{N}$: A directed edge $(i,j)$ between any two vertices $i<j\in\mathbb{N}$ exists independently with probability $p\in(0,1)$. We denote this law by $\mathbb{P}_{p}$. We say that there is a monotone path from $1$ to $n$, write $1\nearrow n$, if there exists a path $1=i_{1}\to i_{2}\to\cdots\to i_{k}=n$ in $G_{p}$ whith $i_{1}<i_{2}<\cdots<i_{k}$. We establish a phase transition for the increasing event
@@
Erdős-Rényi graphs on numbered vertices $\{1,\ldots,n\}$ appeared as an ingredient to recover point configurations in [3]. They prove a sharp threshold at $p=\frac{\log n}{n}$ for the event $1\nearrow n$. Directed acyclic Erdös-Rényi graphs were introduced in [2] and studied as Barak-Erdős graphs since. The articles [8], [9] and [7] established maximal path lengths in Barak-Erdős graphs, both for fixed $p\in(0,1)$ and $p\to 0$.
Numbered vertices can be interpreted as time stamps: each vertex becomes available from its assigned time onward. Related models include random temporal graphs, where monotonicity of paths is considered along edges instead of vertices, see for instance [1], [5] and [4]. These works focus on the length of monotone paths and identify several transition phenomena around $p=c\frac{\log n}{n}$ for both longest and shortest such paths.
A vertex order appears naturally if the nodes correspond to events unfolding in time, as for example in graphical models of causality [10]. Erdős-Rényi graphs directed according to a random vertex order are of particular interest to research in causal discovery [11], where they are widely used in the benchmarking of algorithms that learn causal graphs from data.
Acknowledgments
The first and fourth authors are supported by the ANR grant BISCOTTE - ANR-19-CHIA-0021. The second and third authors are supported by the SuPerGRandMa, the ERG Consolidator no 101087572.
@@ 2 Proof
The central ingredient to prove Theorem 1.1 is the introduction of an exploration process. It indicates the position of having passed a certain number of vertices that are reachable from $1$ by a monotone path. We show that, after suitable centering and scaling, this exploration process converges to a Gumbel variable with a deterministic drift in Section 2.1. This allows to compute explicitly the probability that a monotone path $1\nearrow n$ exists.
@@
the number of vertices up to $k$ that are reachable from $1$ by a monotone path and set $\mathcal{F}_{k}=\sigma(N_{1},\ldots,N_{k})$. A new vertex $k$ is reachable from $1$ precisely when there is an edge to at least one of the $N_{k-1}$ previously reachable vertices. Hence
@@
where each large bullet marks a vertex reachable from $1$ and each group of small dots represents the vertices in the corresponding gap: The $i$-th gap is a geometric variable with parameter $1-(1-p)^{i}$, and the gaps are pairwise independent.
@@ 2.1 Exploration process
denotes the position of the $\lfloor a/p\rfloor$-th vertex reachable from $1$ by a monotone path. Since $P_{p}(a)$ is obtained by summing the successive gap lengths, we have
@@
Define $q_{i}\coloneq 1-(1-p)^{i}$. Assume without loss of generality that $p<1/2$. For any $p<1/2$ and $i\leq\lfloor a/p\rfloor$ it holds $q_{i}\leq 1-(1-p)^{\lfloor a/p\rfloor}\leq B(a)<1$ for some constant $B(a)$ independent of $p$. Each $P_{p}(a)$ is the sum of independent geometric variables with expectation $1/q_{i}$.
@@
so that $Y_{i}$ are exponential variables with intensity $1$ and $X_{i}$ are geometric variables with mean $1/q_{i}$. Then $\log\left(\frac{1}{1-q_{i}}\right)(X_{i}-1)\leq Y_{i}$ which allows the bound
@@
We now conclude the convergence as a process in $a>0$ via a probabilistic version of Dini’s lemma, see [6, Theorem 11.6]: The process
@@
It remains to prove Lemma 2.2.
Proof of Lemma 2.2.
@@ 2.2 Critical window
Proof of Theorem 1.1.
For each $n\in\mathbb{N}$ define the random variable $\mathfrak{X}_{n}$ as the number of vertices smaller than $n$ that are reachable from $1$ by a monotone path, scaled by $p_{n}$, that is